OpenStudy (anonymous):

An artifact was found and tested for its carbon-14 content. If 75% of the original carbon-14 was still present, what is its probable age (to the nearest 100 years)? Use that carbon-14 has a half-life of 5,730 years.

OpenStudy (wolf1728):

elapsed time = half-life*log[(begn amt)/(ending amt)]/log(2) elapsed time = 5,730*log(100/75)/0.3010299957 elapsed time = 5,730*0.1249387366/0.3010299957 elapsed time = 2,378 years